A sociologist finds that in a random sample of 142 U.S. teenagers, 112 regularly use social media to communicate with their friends and family. Use the sample to construct a 90% confidence interval for the proportion of teenagers who regularly use social media to communicate with their friends and family.

What is the margin of error for the 90% confidence interval? Round to three decimal places.
Answer #1

\hat{p}=\frac{x}{n}=\frac{ 112 }{ 142 }= 0.7887

E= Z*\sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}= 1.645 *\sqrt{\frac{ 0.7887 * 0.2113 }{ 142 }}= 0.0564